OpenStudy (anonymous):

In ∆ABC, if a = 8.75 centimeters, c = 4.26 centimeters, and m B = 87°, what is the length of b to two decimal places?

OpenStudy (anonymous):

A)8.92 CM B)9.41CM C)9.53CM D)13.01 E)15.36

OpenStudy (itsyacoolgrl):

you will need the law of cosines to solve it b2=a2+c2−2a×c×cos(B) just substitute your values

OpenStudy (itsyacoolgrl):

Solution: cosB= (a^2+c^2-b^2)/2ac b = sqrt(a^2+c^2-2ac*cosB) = sqrt(8.75^2+4.26^2-2*8.75*4.26*cos87) =9.53 So b=9.53cm

OpenStudy (anonymous):

my teacher hasn't taught us that yet and shes giving us stuff to do can you walk me through

OpenStudy (itsyacoolgrl):

Really? how does she expect you to solve it if you don't know "sohcahtoa" What grade are you in?

OpenStudy (anonymous):

I don't know , ima a sophomore

OpenStudy (itsyacoolgrl):

Hmmm, sorry i'm a freshman. Is this for geometry?

OpenStudy (anonymous):

yes

OpenStudy (itsyacoolgrl):

That's weird :/ I learned sohcahtoa in my freshman year, (I'm going to be a sophomore in the upcoming year). I don't know how to help you but goodluck!